Grade 12th passAlgebra

Find
-3/8 – 7/11
Find
-3/8 – 7/11
Find
-3/8 – 7/11
Find
-3/8 – 7/11

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Take the LCM of the denominators of the given rational numbers.
LCM of 8 and 11 is 88
Express each of the given rational numbers with the above LCM as the common denominator.
Now,
(-3/8)= [(-3×11)/ (8×11)] = (-33/88)
(7/11)= [(7×8)/ (11×8)] = (56/88)
Then,
= (-33/88) - (56/88) ... [∵denominator is same in both the rational numbers]
= (-33 - 56)/88
= (-89/88)

Thanks